OpenStudy (anonymous):

A relationship between x and y is defined by the equation y = -4/3 + 1/3, where x is the input and y is the output. Which statement(s) about the relationship are true? Select EACH correct answer. A. y is a function of x B. The graph of the relationship is a line C. When the input is -3, the output is 4. D. When the input is -2, the output is 3.

OpenStudy (anonymous):

Is the given equation \[y=-4/3+1/3x\]

OpenStudy (anonymous):

oh! no sorry! its y = -4/3x + 1/3

OpenStudy (anonymous):

\[y=(-4/3)x+1/3\] or \[y=\frac{-4}{3x} + 1/3\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

ok :). the last to MC options are easy to check... just sub in the values for x.

OpenStudy (anonymous):

for MC option one, does it have the form of a line? i'e. can you write it as y=mx+b where m,b are numbers?

OpenStudy (anonymous):

I think there are three true statements.

OpenStudy (anonymous):

sorry my internet is being weird!

OpenStudy (anonymous):

yeah i think there's more than one too

OpenStudy (anonymous):

i still dont really get it though....

OpenStudy (anonymous):

ok, lets consider on choice at a time... Choice 1: is y a function of x?

OpenStudy (anonymous):

i dont think so?

OpenStudy (anonymous):

To be a function, each choice of input x, has to have a unique output. i.e. it needs to pass the vertical line test.... Lets two choice two, it will help us decide

OpenStudy (anonymous):

MC 2: is it a line?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

To check, can you write is in the form of a line, y=mx+b? where m,b are just numbers? YES m=-4/3, b=1/3

OpenStudy (anonymous):

Do you see it?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what about the last two?

OpenStudy (anonymous):

ok, back to the first one! 

OpenStudy (anonymous):

does any line pass the vertical line test? i.e. does a vertical line cross our line twice anywhere?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

So it passes, So it's a function!

OpenStudy (anonymous):

nope to there not being a line anywhere**

OpenStudy (anonymous):

so are A and B both answers...?

OpenStudy (anonymous):

k, its a function and it's a line... now does \[(-4/3)(-3)+1/3\] equal 4 like they claim?

OpenStudy (anonymous):

yep, it asks for every true statment(s)

OpenStudy (anonymous):

It equals negative four. I think-

OpenStudy (anonymous):

4+1/3=13/3

OpenStudy (anonymous):

-4*-3=12

OpenStudy (anonymous):

12/3=4

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so it's not c. i'll leave D you you to check.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

D is true.....correct?